OpenStudy (anonymous):

Can some check this my answer The point (6, 8) lies on the line 8x + 11y = 136. Determine if this point is a solution of the system 8x + 11y = 136 0x – y = -8

OpenStudy (anonymous):

8x+11y=136 8(6)+11(8)=136 48+88=136 136=136 true 0x-y=-8 0(6)-(8)=-8

OpenStudy (angela210793):

Yep :)

OpenStudy (anonymous):

ty